# mFabric: An Efficient and Scalable Fabric for Mixture-of-Experts Training

### v1

@@ Abstract.
Mixture-of-Expert (MoE) models outperform conventional models by selectively activating different subnets, named experts, on a per-token basis. This gated computation generates dynamic communications that cannot be determined beforehand, challenging the existing GPU interconnects that remain static during the distributed training process. In this paper, we advocate for a first-of-its-kind system, called mFabric, that unlocks topology reconfiguration during distributed MoE training. Towards this vision, we first perform a production measurement study and show that the MoE dynamic communication pattern has strong locality, alleviating the requirement of global reconfiguration. Based on this, we design and implement a regionally reconfigurable high-bandwidth domain on top of existing electrical interconnects using optical circuit switching (OCS), achieving scalability while maintaining rapid adaptability. We have built a fully functional mFabric prototype with commodity hardware and a customized collective communication runtime that trains state-of-the-art MoE models with in-training topology reconfiguration across 32 A100 GPUs. Large-scale packet-level simulations show that mFabric delivers comparable performance as the non-blocking fat-tree fabric while boosting the training cost efficiency (e.g., performance per dollar) of four representative MoE models by 1.2$\times$–1.5$\times$ and 1.9$\times$–2.3$\times$ at 100 Gbps and 400 Gbps link bandwidths, respectively.
@@
submissionid: #1
@@ 1. Introduction
Mixture-of-Experts (MoE) models (Shazeer et al., 2017; Fedus et al., 2022; mix, [n. d.]c; Zhu et al., 2024; qwe, [n. d.]; gro, [n. d.]; xve, [n. d.]; DeepSeek-AI et al., 2024; dee, [n. d.]) have gained significant traction in the machine learning community to improve the performance of large language models (LLMs) (gro, [n. d.]; mix, [n. d.]c). Unlike traditional methods that scale LLMs by stacking dense layers, which leads to a linear increase in computational costs as model sizes expand, MoE models utilize multiple parallel expert layers and activate only a subset of them based on the input token for each training iteration (e.g., xAI discloses that 25% weights are active in Grok-1 (gro, [n. d.])). This dynamic approach enables models to grow to large sizes without a proportional cost increase in computation.
@@
Today’s GPU interconnects contain intra-server scale-up networks (e.g., NVSwitch (nvs, [n. d.]) or NVLink (nvl, [n. d.]b)) and inter-server scale-out networks (e.g., Ethernet or Infiniband). Both of them are currently dimensioned with uniform and static network topologies (e.g., fully-connected crossbar topology for scale-up networks (nvs, [n. d.]), and Clos-style fat-tree for scale-out networks (Al-Fares et al., 2008; Singh et al., 2015)). To accommodate the temporal and spatial variations of MoE communication patterns, these fabrics contain over-provisioned full bisection bandwidth that is mostly under-utilized. Some recent proposals on the use of optical circuit switching (OCS) to enable topology reconfiguration for spatially non-uniform traffic distributions come with the prerequisite of stable temporal patterns such that no reconfigurations happen during the entire training process (Wang et al., 2023; Jouppi et al., 2023). As a result, these interconnect architectures face bottlenecks, leading to inefficient resource usage and slowdowns in distributed MoE training with both spatial and temporal fluctuations.
Therefore, to fully unlock the computational advantages of MoE models, we need to design a novel GPU interconnect fabric that is adaptable to dynamic all-to-all communication patterns at runtime. This means that the topology reconfiguration needs to happen during the distributed MoE training process. This is very challenging because today’s OCS technologies face fundamental trade-offs between low reconfiguration latency (to enable reconfiguration during training) and high scalability (to interconnect tens of thousands of GPUs) (more details in Table 2).
To understand the problem space, we first perform a comprehensive measurement study in a production cluster to investigate the real-world communication patterns of distributed MoE training. Our measurements reveal that although EP generates strong variability during training, its dynamic range is strictly within an MoE block, creating strong locality for all-to-all traffic on a global scale (§3).
@@
(b) Example of a hybrid parallelism for distributed MoE training that combines DP (gate), EP (parallel expert layers), PP (MoE blocks) and TP (a single expert layer).
Figure 1. Illustration of the MoE model architecture and its distributed training strategies.
Based on this insight, we introduce mFabric, a novel system designed to overcome these challenges by enabling efficient topology reconfiguration during distributed MoE training. At the core of mFabric is a regionally reconfigurable high-bandwidth domain architecture that takes advantage of millisecond-scale OCS for inter-host networks on top of existing electrical interconnects. This hybrid design allows for rapid localized reconfiguration, improving both the scalability and adaptability of the interconnect (§4).
mFabric contains the following key components: 1) mFabric leverages the semi-predictability of all-to-all transfers in the MoE layer to implement a traffic monitor to track regional network demands (§5.1); 2) Based on the obtained demands, mFabric uses a greedy algorithm to generate the OCS topology, balancing algorithmic complexity with topology efficiency (§5.2); 3) Finally, mFabric implements a custom collective communication manager for inter-host DP and EP to route traffics via both electrical interconnects and OCS (§5.3).
To evaluate mFabric, we build a fully functional prototype with 32 NVIDIA A100 GPUs, 16 Mellanox 100G NICs (100, [n. d.]b), a Polatis millisecond-scale OCS (pol, [n. d.]), and a 100G Ethernet switch. In addition, we develop a custom collective communication runtime based on NCCL (ncc, [n. d.]) to support in-training topology reconfigurations. We successfully train three state-of-the-art MoE models using this prototype (§6).
To study the performance of mFabric at scale, we conduct packet-level simulations using four representative real-world MoE models. Our simulation results reveal that mFabric outperforms the state-of-the-art benchmarks, exhibiting comparable training speed to the best electrical (rail-optimized (rai, [n. d.]) and Fat-tree (Al-Fares et al., 2008)) fabrics, while improving cost-efficiency by up to 1.5$\times$ and 2.3$\times$ for 100Gbps and 400Gbps links, respectively. (§7). We also observe that mFabric outperforms TopoOpt (Wang et al., 2023) by up to 2.5$\times$ and exhibits scalability with the cluster size increasing to 30K+ GPUs.
As the first system to enable in-training topology reconfiguration for distributed MoE training, our code will be available at the time of publication to facilitate further discussions.
@@ 2. Background
In this section, we first describe MoE’s model architecture and training parallelization strategies (§2.1). Then, we discuss several existing technologies for GPU interconnects (§2.2).
@@ 2.1. Distributed MoE Training
MoE model architecture. MoE models contain several sequential MoE blocks (layers) (Shazeer et al., 2017; Fedus et al., 2022). As shown in Figure 1(a), each MoE block has one attention layer, a gate unit, and several parallel Feed-Forward Networks (FFNs) called experts. The input token $x$ is first fed into the attention layer. After that, the gate unit to select the most relevant experts based on the output of the attention layer. For example, expert 2 is activated while other experts are omitted. This is called computation-based routing and is the key to enabling MoE’s sparse architecture that scales to trillions of parameters without a linear increase in computation cost. The output token, $y$, is the weighted sum of the outputs of all selected experts.

### v2

@@ Abstract.
Mixture-of-Expert (MoE) models outperform conventional models by selectively activating different subnets, named experts, on a per-token basis. This gated computation generates dynamic communications that cannot be determined beforehand, challenging the existing GPU interconnects that remain static during distributed training. In this paper, we advocate for a first-of-its-kind system, called mFabric, that unlocks topology reconfiguration during distributed MoE training. Towards this vision, we first perform a production measurement study and show that the MoE dynamic communication pattern has strong locality, alleviating the need for global reconfiguration. Based on this, we design and implement a regionally reconfigurable high-bandwidth domain that augments existing electrical interconnects using optical circuit switching (OCS), achieving scalability while maintaining rapid adaptability. We build a fully functional mFabric prototype with commodity hardware and a customized collective communication runtime. Our prototype trains state-of-the-art MoE models with in-training topology reconfiguration across 32 A100 GPUs. Large-scale packet-level simulations show that mFabric achieves performance comparable to a non-blocking fat-tree fabric while boosting the networking cost efficiency (e.g., performance per dollar) of four representative MoE models by 1.2$\times$–1.5$\times$ and 1.9$\times$–2.3$\times$ at 100 Gbps and 400 Gbps link bandwidths, respectively.
@@
submissionid: #1conference: SIGCOMM’25; September 8 - 11, 2025; Coimbra, Portugal
@@ 1. Introduction
Mixture-of-Experts (MoE) models (Shazeer et al., 2017; Fedus et al., 2022; mix, [n. d.]c; Zhu et al., 2024; qwe, [n. d.]; gro, [n. d.]a; xve, [n. d.]; DeepSeek-AI et al., 2024; dee, [n. d.]; DeepSeek-AI et al., 2025; Team, 2024) have gained significant traction in the machine learning community to improve the performance of large language models (LLMs) (gro, [n. d.]a; mix, [n. d.]c). Unlike traditional methods that scale LLMs by stacking dense layers, which leads to a linear increase in computational costs as model sizes expand, MoE models utilize multiple parallel expert layers and activate only a subset of them based on the input token for each training iteration (e.g., xAI discloses that 25% weights are active in Grok-1 (gro, [n. d.]a), while DeepSeek-V3 (dee, [n. d.]) only activates 37 billion parameters over total 671 billion parameters). This dynamic approach enables models to grow to large sizes without a proportional cost increase in computation.
@@
Today’s GPU interconnects contain intra-server scale-up networks (e.g., NVSwitch (nvs, [n. d.]) or NVLink (nvl, [n. d.]b)) and inter-server scale-out networks (e.g., Ethernet or Infiniband). Both of them are currently dimensioned with uniform and static network topologies (e.g., fully-connected crossbar topology for scale-up networks (nvs, [n. d.]), and Clos-style fat-tree for scale-out networks (Al-Fares et al., 2008; Singh et al., 2015)). When accommodating the temporal and spatial variations of MoE communication patterns, these fabrics contain over-provisioned full bisection bandwidth that is mostly under-utilized. Some recent proposals on the use of optical circuit switching (OCS) perform topology reconfiguration for spatially non-uniform traffic distributions. However, they assume stable temporal patterns such that no reconfigurations occur during the entire training process (Wang et al., 2023; Jouppi et al., 2023). As a result, these interconnect architectures face bottlenecks, leading to inefficient resource usage and slowdowns in distributed MoE training.
Therefore, to fully unlock the computational advantages of MoE models, we need to design a novel GPU interconnect fabric that is adaptable to dynamic all-to-all communication patterns at runtime. Achieving such adaptability requires the topology to be reconfigurable during the distributed MoE training process. This is very challenging because today’s commodity OCS technologies face fundamental trade-offs between low reconfiguration latency (to enable reconfiguration during training) and high scalability (to interconnect tens of thousands of GPUs) (more details in Table 1).
To understand the problem space, we first perform a comprehensive measurement study in a production GPU cluster to investigate the real-world communication patterns of distributed MoE training. Our measurements reveal that although EP generates substantial variability during training, its dynamic range is strictly within an MoE block, creating strong locality for all-to-all traffic on a global scale (§3).
@@
(b) Example of a hybrid parallelism that combines DP (gate), EP (parallel expert layers), PP (MoE blocks) and TP (a single expert layer).
Figure 1. Illustration of the MoE’s gated expert architecture and its distributed training strategies.
Based on this insight, we introduce mFabric, a novel system designed to overcome these challenges by enabling efficient topology reconfiguration during distributed MoE training. At the core of mFabric is a regionally reconfigurable high-bandwidth domain based on millisecond-scale reconfigurable OCS that sits at the boundary between scale-up and scale-out networks. This design augments the existing static electrical interconnect with the capability of rapid regional reconfiguration while still preserving its scalability (§4).
mFabric contains the following key components: 1) mFabric exploits the partially predictable nature of all-to-all communications to track regional traffic demands at runtime (§5.1); 2) Based on the obtained demands, mFabric uses a greedy algorithm that generates the tailored network topology and reconfigures the OCS to realize it (§5.2); 3) With the reconfigured topology, mFabric uses a customized collective communication runtime to orchestrate inter-host DP and EP communications on EPS and regional OCS fabrics (§5.3).
To demonstrate mFabric, we build a fully functional prototype with 32 Nvidia A100 GPUs, 16 Mellanox NICs (100, [n. d.]b), a Polatis millisecond OCS (pol, [n. d.]), and an Ethernet switch. In addition, we develop a custom collective communication runtime based on Nvidia collective communications library (NCCL) (ncc, [n. d.]) to support in-training topology reconfigurations. Using this prototype, we successfully demonstrated the benefits of mFabric on state-of-the-art MoE models (§6).
To evaluate the performance of mFabric at scale, we perform packet-level simulations using four representative real-world MoE models. Our results reveal that mFabric outperforms the state-of-the-art fabrics, exhibiting a training speed comparable to Rail-optimized (rai, [n. d.]) and Fat-tree (Al-Fares et al., 2008) fabrics while significantly improving cost-efficiency. Specifically, mFabric improves networking infrastructure cost-efficiency by 1.2$\times$–1.5$\times$ (1.9$\times$–2.3$\times$) compared to Fat-tree and 1.4$\times$–1.5$\times$ (2.3$\times$–2.4$\times$) compared to Rail-optimized for 100Gbps (400Gbps) links. We also observe that mFabric outperforms TopoOpt (Wang et al., 2023) by up to 2.5$\times$ and supports scalability with the cluster size increasing to 30K+ GPUs (§7). When connected to co-packaged optical ports attached to GPUs, mFabric augments high-radix scale-up systems like NVL72 by 1.3$\times$8).
As the first system to enable in-training topology reconfiguration for distributed MoE training, we will release our code by the time of publication.
@@ 2. Background
In this section, we first describe MoE’s model architecture and training parallelization strategies (§2.1). Then, we discuss several GPU interconnects for distributed training (§2.2).
@@ 2.1. Distributed MoE Training
MoE model architecture. MoE models contain several sequential MoE blocks (Shazeer et al., 2017; Fedus et al., 2022). As shown in Figure 1(a), each MoE block has an attention layer, a gate unit, and several parallel feed-forward networks (FFNs) called experts. The input token $x$ is first fed into the attention layer. After that, the gate unit selects the most relevant experts based on the output of the attention layer. This is called computation-based routing and is the key to enabling MoE’s sparse architecture that scales model parameters without a linear increase in computation cost. The output token, $y$, is the weighted sum of the outputs of all activated experts.